OpenStudy (anonymous):

A 12-ounce soda can is 5 inches tall and 2.5 inches across. How much soda can it hold if it is completely full? Round your answer to the nearest tenth of a cubic inch.

OpenStudy (chealyn98):

V = pi * r^2 * h V = 3.14 * 1.25^2 * 5 V = 24.5312 = 24.5 cubic inches